Question 1 Use Adam-Bashforth-Moulton two-step explicit and implicit methods to approximate y(2.4) for the following differential equation with y(2)=14.7781 and y(2.2)=19.855 USE FOUR DECIMAL DIGIT ROUNDING.-y-y/x=0
Answer #1

Solution :- Given that y (2, 4) Differential eanation with y (2) = 14.7781 (2.2) -19-855 f(x,y) = y + Ya het lo=2 , y = 14.77